### v5

# SPA: Towards A Computational Friendly Cloud-Base and On-Devices Collaboration Seq2seq Personalized Generation

Abstract
Large language models(LLMs) have shown its outperforming ability on various tasks and question answering. However, LLMs require substantial memory storage on low-resource devices. More critically, the computational speed on these devices is also severely limited. In this paper, we propose SPA(Side Plugin Adaption), a lightweight architecture for fast on-devices inference on the constraints of strict on-devices computation and memory constraints. Compared with other on-devices seq2seq generation, SPA could make a fast and stable inference on low-resource constraints, allowing it to obtain cost effiency. Our method establish an interaction between a pretrained LLMs on-cloud and additive parameters on-devices, which could provide the knowledge on both pretrained LLMs and featured personal feature. Further more, SPA provides a framework to keep feature-base parameters on low computational devices while leave the parameters containing general information on the high computational devices.  [^1] Our code is public at https://github.com/OceannTwT/spa
1 Introduction
Large Language Models (LLMs)[1, 2, 3] have demonstrated incredibly powerful capabilities in supporting downstream tasks. These models perform exceptionally well across various downstream tasks. Additionally, different LLMs such as GPT-4[4], Llama[3], and ChatGLM[5, 6] exhibit similar abilities in handling tasks across multiple dimensions[7]. However, deploying and computing LLMs on resource-constrained devices introduces significant memory and computational pressure on these devices[8]. Therefore, researching reliable methods for deploying LLMs on-devices becomes critically urgent.
Figure 1: Three methods for model deployment are based on different combinations of cloud and device setups.
To address the issue of deploying large language models (LLMs) on devices, various methods have been proposed to adapt them for edge devices. Subformer [9] reduces storage and computation requirements by sharing parameters. Similarly, EdgeFormer [10] proposes a balanced sharing scheme, exploring the impact of different architectures on device deployment. LLMCad [11] employs a collaborative model configuration to achieve lightweight deployment, where a small model generates simple words and a high-accuracy model verifies and corrects errors.
Additionally, research has proposed parameter-efficient fine-tuning methods for LLMs [12, 13, 14], utilizing adapter strategies to flexibly adapt to downstream tasks while preserving model capabilities. Adapters are stored separately from the base model and require only a few parameters to customize for downstream tasks [15]. For on-device deployment, adapters can be used for incremental deployment, integrating incremental knowledge into existing models to support downstream tasks and obtain user attributes [16]. However, this approach requires consideration of data transmission requirements between devices [17], and reducing transmission latency is a significant challenge in the design of separated models.
Our research proposes a method called Side Plugin Adaptation (SPA), which achieves the deployment of large models on-devices by separating adapters from pre-trained models. Specifically, our method introduces an side adapter classifier. This classifier allows us to choose between leveraging the inherent capabilities of the original model and integrating feature information generated by adapters. In this way, we aim to minimize the frequency of adapter usage while maximizing the utilization of common-sense knowledge inherent in the original large language model. This strategy offers a dual advantage: leveraging the unique feature attributes of adapters while maximizing the inherent capabilities of a widely based model. Compared to traditional on-devices model setups, our method significantly improves inference speed.
We comprehensively evaluate our approach on various datasets across different domains. These evaluations not only emphasize the advantages of our method compared to traditional approaches but also highlight its potential in improving overall model capability while achieving significant adapter effects. In addition to enhancing text generation and semantic understanding services, our approach demonstrates a stronger ability to understand textual features, better meeting user needs.
Our contributions can be summarized as follows:
  • We innovatively propose the SPA framework, deploying large models on-device via adapters. Our method significantly reduces the difficulty of fully deploying large models on the edge and presents a feasible and efficient solution.
  • We introduce a classifier to determine whether to use adapter content for inference. This innovative classifier design significantly improves the efficiency of cloud server to edge transmission while maintaining task performance.
  • SPA has been thoroughly evaluated across multiple datasets. Experiments show that our framework performs well on-devices. Additionally, SPA can achieve results close to the original model with lower transmission latency.
2 Related Work
2.1 Parameter-Efficient Transfer Learning
In recent years, efficient parameter-efficient transfer learning has become a popular method for fine-tuning large language models. It achieves performance close to full fine-tuning by refining a subset of parameters. Among these methods, adapters, as introduced by [12], incorporate two additional feed-forward layers, enhancing their adaptability to downstream tasks. Similarly, LoRA [13] employs a rank reduction followed by rank promotion method on the original language model. In the context of language model fine-tuning, prefix tuning [18, 19] involves introducing a trainable prefix vector before the keys and values in the multi-head attention mechanism, producing effects similar to the first two methods.
The MAM adapter [20] synthesizes these techniques and identifies superior results in parameter fine-tuning. Moreover, LST [21] provides a method for saving space during parameter fine-tuning. It achieves optimal parameter fine-tuning with the same memory consumption by leveraging adapters on the device. This approach significantly reduces memory usage when training large language models on edge devices. Meanwhile, the QLora[22] method combines quantization rules with LoRA’s rank promotion and demotion scheme, optimizing training resources and space savings.
Furthermore, the CuQA[14] method offers a classifier that selects between the original language model and edge models based on the relevance of training and testing data during training. Through this selection process, we observe that efficient parameter fine-tuning methods can effectively embed new knowledge, thereby expanding existing language models to incorporate new knowledge bases in support of specific tasks. Our method leverages the parameter-saving capabilities of efficient parameter-efficient transfer learning and the adaptive information contained in parameter fine-tuning inspired by CuQA. We provide robust support for various generative tasks, demonstrating enhanced capabilities.
2.2 On-device ML optimization
For on-device sequence-to-sequence models, various techniques have been adopted to reduce the parameters of pre-trained models, including methods such as quantization[23, 24] to reduce model size, knowledge distillation[25, 26] techniques that learn from large models and transfer knowledge to smaller models, and techniques such as compression[27, 28] and pruning[29, 30, 31] to reduce the number of model parameters. Some methods like Edgeformer [32] achieve model reduction by sharing model parameters. Low-rank decomposition is also used to reduce model size [33]. Methods like Delight [34] achieve model size reduction by compressing attention mechanisms. These techniques are adjusted and adapted based on the parameters of the Transformer [35] architecture, and they yield effective parameter reduction results for models with relatively fewer parameters.
Given that a considerable portion of these methods can only be applied to models with small model, they may not produce substantial optimizations in the context of large models[36]. Computational and storage resources on devices may not be sufficient to accommodate LLMs. Therefore, we propose a framework for hybrid cloud server and on-device sequence-to-sequence model. The framework aims to meet the needs of certain downstream tasks and user characteristics by combining the capabilities of the cloud and edge.
3 Methodology
Considering the edge computing and storage capabilities, as well as the need to learn user characteristics and contextual information related to text generation, we propose SPA, a lightweight architecture for fast on-devices inference on the constraints of strict on-devices computation and memory constraints. With this design, we can evaluate its support for downstream tasks and compare it with the results of the original edge adapter.
3.1 Side Model Design
Collaborative learning and training between edge and cloud models currently faces the following major bottlenecks[37]: While all models are deployed on edge devices, the computation required for the models involves extensive floating-point operations, and the memory needed for model training significantly increases[38]. These issues make it tough to directly use LLMs for inference on edge devices. On the other hand, when data is placed in the cloud, although general computing power is ensured, some personalized features cannot be efficiently transferred from the edge to the cloud for effective learning. These bottlenecks highlight the urgent need to design a model that combines both edge and cloud capabilities to better adapt to the respective requirements.
We consider placing some learnable parameters on the side to enable rapid personalization of information learning settings. In general, SPA involves using a classifier to select the output results generated by the original LLM or the side block. The classifier chooses the path that aligns better with the desired features, thereby optimizing the quality of text generation. Additionally, the classifier can decide whether to combine the edge model with the base model stored in the cloud. In this way, the edge model only needs to learn the feature differences relevant to the current task, while the general information is provided by the base model, reducing the resource latency costs on edge devices.
Specifically, The function $f$ is the single layer on origin LLMs, and the function $g$ is the corresponding layer on side model. the methods above could conclude to these two functions:
$\mathcal{H}_{i}(x)=f_{i}(x)+\sigma_{i}g_{i}(x)$ (1)
We employed a linear layer of $H\times 2$ as our classifier. The advantage of this design lies in our ability to use the Transformer’s output as the hidden state. This hidden state is then softmaxed to a range between $[0,1]$, thereby obtaining a specific distribution for classification choices $\sigma_{i}=\arg\max{\textbf{M}_{c}f_{i}(x)}$.
Through such a linear mapping, the model can learn whether predicting a certain token is better accomplished by selecting the existing model’s output. This approach typically leverages previously predicted information like: Los Angeles, Which Los is predicted in previous token. Alternatively, the model can employ a side-end incremental model that relies more on user-specific feature information. This helps in determining the optimal choice. Since $\sigma_{i}$ is learnable, the loss function for the whole model can be depicted as follows:
Figure 2: Cloud and Device collaboration model SPA. This framework utilizes a classifier to provide optional downstream task computation choices for edge devices, thereby reducing computational pressure on edge devices and improving the overall performance of specialized device models.
$\mathcal{L}_{LM}=-\sum_{i=1}^{T}\log_{(\Theta+\sigma\Delta\Theta)}{\textbf{P}(x_{i}|x_{<i})}$ (2)
Here, we keep $f(x)$ frozen and unchanged, while training the side model $g(x)$ and the classifier $\sigma$. Unlike the fixed classifier $\sigma^{\prime}$ used for text similarity in the original model, our classifier $\sigma$ based on a linear layer can determine whether the predicted word exists in the knowledge base of the underlying language model. This approach reduces the inference time for highly repetitive word sequences, stores more general information in the cloud, and generates user-specific features on edge devices. Compared to EdgeFormer, the main bottleneck of our model is the communication latency between the edge device and the server [17]. This can be optimized by controlling the frequency and timing of communication. EdgeFormer [32] achieves a smaller model with similar functionality but fewer FLOPs by reducing the original model’s size (e.g., by four or eight times) and sharing parameters. However, this design requires higher computational power and specialization for edge devices. While our approach results in a smaller edge model, its prediction speed relative to the total latency is very fast. In this case, the computational power of the edge device will not become a bottleneck for the prediction results.
3.2 Communications Latency
For communication between cloud servers and edge models, we adopt the architecture used in LST. LST improves memory capacity through backpropagation. For prediction, if our transformer has $L$ layers, the Parallel adapter or LoRA method results in higher communication latency costs. Specifically, the latency includes three parts: $T_{pretrained}$ represents the computation latency of the original pre-trained model, $T_{on-devices}$ represents the computation latency of the edge device model, which could be the Adapter part, LoRA part, or LST part used for edge device inference. Additionally, $T_{net}$ is the network latency when transmitting data between the cloud and devices. Then, the total latency can be evaluated as:
$T_{total}=T_{on-devices}+T_{pretrained}+T_{net}$ (3)
While $T_{pretrained}$ keep the same for identical model. For model on edge devices, the inference latency is lay on the CPU or GPU computing capability, denoted as $f^{e}$. $F_{data}$ is the FLOPs required to compute on edges, $C_{devices}$ indicates the numbers of CPU or GPU are available to compute on this model. Then:
$T_{on-devices}=\frac{F_{data}\times{C_{devices}}}{f^{e}}$ (4)
For the network latency when transmitting data between cloud and devices. Assume the TCP/IP connection latency is $\tau$, The $T_{net}$ could be formulated as:
$T_{net}=\mathcal{M}\times(\tau+T_{data})$ (5)
Where $\mathcal{M}$ is the transmission times for one token prediction, it would be changed for different architecture, because we have to transfer back the hidden state from side model to base model. SPA effectively merges the process in scenarios where edge-cloud connectivity is involved.
4 Experiments Setup
4.1 Datasets
We evaluated our approach on the following training dataset.
  1. 1.
    XSum[39]: There are 227, 000 summaries of online articles generated by the BBC available on XSum, and we assessed the relationships between the articles and their corresponding summaries.
  2. 2.
    CNN Daily Mail[40]: The CNN Daily Mail dataset consists of 312, 000 summaries of CNN Daily Mail articles and its origin articles.
  3. 3.
    CoQA[41]: CoQA datasets contains 127, 000 questions encompassing 8, 000 conversation-based question-answer pairs, constituting a large-scale QA system spanning various domains.
  4. 4.
    SciQ[42]: The SciQ dataset comprises 13, 679 publicly available scientific exam questions covering diverse disciplines such as biology, chemistry, physics, and others. Answers to specific questions are obtained through contextual information.
4.2 Baselines
We compare our approach with various efficient parameter tuning methods, including three baseline: LoRA[13], Adapters[12], and LST[21]. These methods are widely used in different downstream tasks and have achieved outstanding results.
4.3 Models
We utilize Llama[3], an open-source large language model, and employed the PEFT method to evaluate task inference latency, network latency, and task performance. We conducted our analysis using the Llama-7B model. This open-source large language model demonstrates very strong capabilities in general domains.
4.4 Implement Details
[…]
Table 1: Latency evaluated from Different efficient tuning methods. Measuring new parameters percentage rate, Inference latency, Network latency and Total latency on prediction for 50 new tokens. Ratios indicates the ratios of direct transmission times and pretrained model inferencing times. The Rouge-L scores on different PEFL methods and our model on various datasets.
Model % Param Infer. latency Net. latency Total latency Ratios XSum CoQA
LLaMA-7B + LoRA 0.19 3.26 6.37 9.63 32.0 38.24 40.32
LLaMA-7B + Adapter 0.38 3.32 12.56 15.88 64.0 38.18 40.51
LLaMA-7B + LST 0.2 3.29 0.31 3.60 1.0 28.78 31.24
LLaMA-7B + SPA 0.21 3.30 0.18 3.48 0.62 35.52 37.30
5 Experiments
5.1 Main Result
The performance of SPA surpasses that of LST. Table 1 shows that, under the premise of lightweight parameters, the blocks trained by SPA can effectively guide the original model. Consequently, even with a smaller computational load at the edge, the model can exhibit strong inference capabilities and enhance performance for the given task.
In the SPA architecture, network transmission is a crucial aspect, particularly in the context of SPA. Therefore, as the discussion in Section 3, it’s essential to compare the previous model with the SPA model and estimate the corresponding latency required for predicting the entire model. Given that our model is relatively large, we cannot deploy the entire model at the edge as in the case of Edgeformer. In the collaborative process between cloud-based and on-device components, transmitting critical information is necessary. Consequently, we will be conducting inference latency experiments to assess this aspect.
5.2 Latency on Prediction
To address concerns about network latency and overall inference delay, we conducted comparative experiments on data exchange for these models. Additionally, we compared the original Adapter, LoRA, and LST methods, evaluating them based on data exchange times and overall model latency. We performed these assessments using the same training configuration as mentioned above.
We also uniformly evaluated the time taken to generate 50 tokens. Additionally, our single-sided model and the baseline model were placed within a server cluster to simulate network transmission delays in an ideal network environment. This was done to assess the corresponding network transmission latency. The respective metric results obtained from the testing are recorded in Table 1. the difference between single and multiple communication latency is substantial, with inference costs relying mainly on cloud server computing power.
Table 2: The Rouge-L scores on different settings and SPA model on various datasets.
Model Ratios XSum CNN-DM CoQA SciQ
LLaMA-7B + Fully Integration 1.0 36.24 40.55 38.81 25.91
LLaMA-7B + Fully Separation 0.68 30.29 34.19 32.11 23.71
LLaMA-7B + LST 1.0 28.18 32.15 31.24 23.24
LLaMA-7B + SPA 0.64 36.52 39.22 37.30 25.38
The SPA model outperforms other methods in terms of network latency. Notably, SPA has many excellent features. It can make predictions by choosing between the pre-trained model and the integrated adapter model through a classifier. Due to the use of the LST structure, SPA’s architecture is more independent on devices, giving it an advantage over methods like LoRA and Adapters.
5.3 Impact of Side Model
We have assessed the impact of the edge model on the overall model and analyzed the word repetition rates between the edge model and the original model. Based on this analysis, we evaluate whether the results of our model meet our expected outcomes when selecting the edge model for inference. This metric can also be used to assess the extent of the difference between the new model and the original model and the proportionality with the edge model we have chosen. We show the evaluation of side model effect on Table 2.
We could find that while the origin cloud model and on-devices model generate different results, the classifier $\sigma_{i}$ would make a option that choose a more ideal information from the two alternations. When it comes to a same result, the classifier would choose to pick the cloud one for less network communication cost.
5.4 Training Scales Analysis
To assess the impact of training data sizes on model performance, we train the SPA model and utilize different sizes of training sets, as specified below:
We train the models using the same network architecture and hyperparameters, with a consistent training epoch of 15 rounds. Text generation is performed using beam search.
From the results, the model’s performance improves across multiple datasets with an increase in training data sizes, especially from small to medium data. This indicates that augmenting training data enhances the model’s generalization ability. However, as data sizes continues to increase to the full set, the performance improvement becomes less pronounced, indicating the presence of marginal effects. Considering the escalating computational resources, medium data size offers a favorable performance-to-cost ratio.
Table 3: The Rouge-L scores on different training data sizes on various datasets.
Data sizes Xsum CNNDM CoQA SciQ
$\text{SPA}_{\text{small}}$ 31.5 36.2 34.7 22.1
$\text{SPA}_{\text{medium}}$ 33.8 38.4 36.2 23.8
$\text{SPA}_{\text{full}}$ 35.5 39.2 37.3 25.4
Additionally, we record the proportion of adapter usage by the model, as shown in Table 5. As data sizes increases, the percentage of adapter usage decreases, with a higher probability of selecting the original model’s output. This further validates that increasing data sizes enhances the model’s inherent language understanding and generation capabilities. Increasing the size of training data significantly enhances performance on downstream tasks. However, the marginal benefits decrease as data size continues to grow. Considering computational resources, a medium-sized dataset offers the best performance-to-cost ratio. Additionally, more data reduces the need for adapters, thereby improving the model’s language understanding capabilities.
5.5 Alternatives on Side Plugin
SPA outperforms both LST and the pre-trained model, showcasing better expressive capabilities. we conduct ablation experiments for identifying the source of this enhanced performance.
We present three configurations in our study. In the first setup, referred to as Fully Integration. The edge adapter directly interfaces with the original model. This arrangement allows the edge adapter to effectively learn and correct errors within the original model. In addition, the setup of S1 can be viewed as a parallelized LST setup, which has the advantage that he may perform better on generative tasks than the original serial option of LST.
Figure 3: Performance on XSum datasets between LST and SPA.
Figure 4: Performance on CNN-DM datasets between LST and SPA.
In the second setup, a choice is made between the edge-adapted model and the original model. The classifier utilizes the original model’s output to make decisions between the results of the original model and the edge-adapted model. We denote the setting as Fully Separation In this case, the edge on-devices model operates with a higher degree of independence from the original model.
Table 4: The Side Model Effect evaluation.
Model % Param M Diff rates Diff Pick rates
LLaMA-7B + SPA 0.21 0.62 78.2 56.8
The final setup involves the original LST framework, which can be viewed as relying entirely on the edge adapter for result generation. By comparing our model with these three setups, we aim to determine which configuration best aligns with the requirement for consistency between efficiency and capability that the model needs to achieve. The capabilities demonstrated by different experiments are summarized in Table 2.
The performance improvement in downstream tasks mainly depends on the adapter’s adjustments to capture incremental differences. As shown in Table 2, while introducing the classifier in the SPA model causes a slight performance drop compared to integrating adapters in all scenarios, this reduction is minimal. However, the increase in inference speed is significant. Given our need for quick responses at the edge, this classifier aligns well with our vision and requirements. For more fundamental information, predictions are made using the original model. The classifier greatly reduces latency from continuous calls to the edge-side model during predictions, ensuring effective learning of both user and task features.
Parallel processing. We also evaluate the performance of the sequence to sequence task between the original model and the model after parallel processing, and make corresponding comparisons under different datasets, and determine its performance under the corresponding task by the Rouge-L metrics.
In the parallelized design, the performance for our generative class of tasks is consistently higher than the original serialized design. As shown on Figure 4, such a parallelized design also highlights the plug-and-play nature of our model. After a very small number of rounds, the performance of parallelization is already more prominent than 20 rounds of generative training. Thus the use of a parallelized design in this paper is also in line with the corresponding sequence-to-sequence generation task.
Figure 5: The summary case study on classifier option, The pretrained model choices were marked as blue, while SPA model choices were marked as black.
5.6 Case Study
As shown on Figure 5. By detecting and processing the impact of classifiers on test cases, including the labels and branches used, we assess their influence and examine whether predictions are based on the original corpus or newly embedded knowledge. We analyze the top five results, observing key activation moments of the end-to-end model, as well as the predictions of the original pre-trained model at these moments.
Table 5: Evaluations on side model ratios.
Data sizes $\text{SPA}_{\text{small}}$ $\text{SPA}_{\text{medium}}$ $\text{SPA}_{\text{full}}$
Usage Percentage 82.3 71.5 61.8
The original model is primarily based on the usage of existing corpora and generated pronouns. These pronouns are generated within the original LLM, enabling it to learn the desired outcomes that users want to predict. For specific feature information, the output is primarily associated with embedded sentiment words and pronouns. These features are learned from the SPA model and are applicable to user profile features or downstream task requirements.
6 Conclusions
This paper introduces a model suitable for cloud-base and on-devices collaboration. The model leverages the capabilities of LLMs for sequence-to-sequence tasks and utilizes user-specific features from the edge model to guide the overall large model towards our desired objectives. Through a classifier, we can select the most suitable generated sentences. Our approach presents a new paradigm in addressing issues related to edge model deployment, providing valuable insights for future work. We believe that this approach is groundbreaking and can serve as a valuable reference for future research.
Appendix A Limitation
Our model is a cloud-based and on-device cooperative system, which exhibits a strong inference ability on various downstream tasks. Leveraging the high-parameter nature of the large model, it can achieve excellent end-to-end prediction results. However, our model still has some shortcomings. For instance, its collaborative process inevitably introduces some latency. In scenarios that demand high network resources, our model may not perform optimally for reasoning and training.
Additionally, there may be privacy concerns since our corpus is ultimately processed in the cloud, allowing cloud services to access the corpus content. Nonetheless, we protect privacy to a certain extent by keeping the model-related parameters locally. We only need these local parameters to perform inference based on user features. These issues warrant further investigation and improvement in future work.